# Effect of Overheat and Direct Flow Loading on the MEMS Bistable Flow Sensor

### v1

@@ 3.4 Dynamics of the ST transition
In contrast to previous studies kessler2020flow ; kessler2021sampling , where a single-point LDV system was employed to measure only the midpoint deflection, our approach in this study takes a significant step forward. We utilize the Micro System Analyzer (Polytec, MSA 600).
The MSA operates in a scanning mode, allowing us to capture the complete shape of the entire beam during the transient ST and SB events. To achieve this, the MSA records a 36 milliseconds long velocity signal from 56 points distributed along the length of the beam. We actuate the beam buckling with a linear ramp of $V_{\mathrm{ES}}$ ranging from 0 to 200 V at a frequency of 400 Hz. The displacement of the beam at each point is determined by integrating the velocity signal over time. Phase averaging is applied to calculate the average displacement at each measurement point on the beam, effectively eliminating free vibrations following ST and SB.
@@
Figure 11: Snapshots of beam deformation during snap-throught transition (a), and snap-back transition (b).
@@
This represents a real-time beam profile measurement during the dynamic snap-through and snap-back transitions. Notably, we observe a slight asymmetry in the shape of $z(x,t)$ during both transitions concerning the midpoint, $z(x=0.5L)$. This implies that the transitions occur at lower $V_{\mathrm{ST}}$ and higher $V_{\mathrm{SB}}$ values than what would be anticipated from a symmetric first-mode transition. The presentation of these asymmetric snapshots of beam deformation during ST and SB events is a novel contribution, supporting the hypothesis presented in medina2012symmetry regarding the nature of this transition. Additionally, it provides insight into the shape of free vibrations following snap-through, in line with findings from song2021free .
Notably, during the beam transition, it is evident that the center of the beam lags behind the rest of the beam, with the exception of the edges, which remain static. Particularly, the sections between the center and the edges, around $0.40.2<x/L<$ 0.4 and $0.80.6<x/L<$ 0.8, lead the motion during a significant portion of the deformation. This shape likely reflects a combination of various modes and underscores the contribution of the second symmetric mode. This discovery encourages further exploration in this direction.
@@ 4 Summary and conclusions
This study delved into critical aspects of the sensor’s behavior. Firstly, we initiated by quantifying the “loading effect” induced by the laser Doppler vibrometer, shedding light on the impact of laser heating on voltage measurements.
Secondly, we investigated the beam’s response to varying levels of Joule heating. While the heating voltage $V_{\mathrm{h}}$ influenced the absolute values of $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ voltages, the device’s sensitivity remained notably consistent. Our analytical model supported these findings and revealed a latching effect with increasing Joule heating. These results suggest that the beam overheat ratio may not be as pivotal as assumed for most thermal MEMS sensors, promising substantial energy consumption reduction.
Thirdly, we investigated the influence of the flow angle on the performance of the beam. We concluded that the device performs similarly to previous hot-wire-based microsensors and can measure flow direction.
We conducted a detailed analysis of the beam’s response to fluctuations at different frequencies, typically present in various applications. Our results confirmed that this device, operating near critical transition points, is also sensitive to turbulent flow velocity fluctuations.
Finally, we obtained the first direct beam shape measurements during transitions and confirmed the theoretical predictions of asymmetric, higher-than-first-mode transitions. We can confirm that the single degree-of-freedom approximation of the single of beam needs further improvement.
Overall, this study lays the groundwork for designing a series of flow sensors and multi-sensor arrays for velocity and velocity gradient measurements using initially curved bistable microbeams. However, unanswered questions still require further investigation, such as the sensor’s response to out-of-plane flow directions, its performance in realistic environments, and the noise levels at which it can accurately measure low and high flow velocities for different applications.
@@ Acknowledgments
The authors acknowledge the financial support of the Israeli Innovation Authority, NOFAR program and of RAFAEL LTD, Advanced Defense Systems. S. Krylov acknowledges the support of the Henry and Dinah Krongold Chair of Microelectronics. The authors would also like to thank Gal Spaer Milo, Grigori Gulitski, and Stella Lulinsky for their assistance.

### v2

@@ 3.4 Dynamics of the ST transition
In contrast to previous studies [21, 22], where a single-point LDV system was used to measure only the deflection at the midpoint, our approach in this study takes a significant step forward. We utilize the Micro System Analyzer (Polytec, MSA 600).
The MSA operates in scanning mode, allowing us to capture the complete shape of the entire beam during the transient ST and SB events. To achieve this, the MSA records a 36 milliseconds long velocity signal from 56 points distributed along the length of the beam. We actuate the beam buckling with a linear ramp of $V_{\mathrm{ES}}$ ranging from 0 to 200 V at a frequency of 400 Hz. The displacement of the beam at each point is determined by integrating the velocity signal over time. Phase averaging is applied to calculate the average displacement at each beam measurement point, eliminating free vibrations after ST and SB.
@@
Figure 11: Snapshots of beam deformation during snap-through transition (a), and snap-back transition (b).
@@
This represents a real-time beam profile measurement during the dynamic snap-through and snap-back transitions. In particular, we observe a slight asymmetry in the shape of $z(x,t)$ during both transitions concerning the midpoint, $z(x=0.5L)$. This implies that the transitions occur at lower $V_{\mathrm{ST}}$ and higher $V_{\mathrm{SB}}$ values than what would be anticipated from a symmetric first-mode transition. The presentation of these asymmetric snapshots of beam deformation during ST and SB events is a novel contribution, supporting the hypothesis presented in [23] regarding the nature of this transition. Additionally, it provides insight into the shape of free vibrations following snap-through, in line with findings from [39].
Notably, during the beam transition, it is evident that the center of the beam lags behind the rest of the beam, except the edges, which remain static. Notably, the sections between the center and the edges, around $0.40.2<x/L<$ 0.4 and $0.80.6<x/L<$ 0.8, lead the motion during a significant portion of the deformation. This shape likely reflects a combination of various modes and underscores the contribution of the second symmetric mode. This discovery encourages further exploration in this direction.
@@ 4 Summary and conclusions
This study delved into critical aspects of the sensor’s behavior. First, we quantified the “loading effect” induced by the laser Doppler vibrometer, which sheds light on the impact of laser heating on voltage measurements.
Second, we investigated the beam’s response to varying levels of Joule heating. While the heating voltage $V_{\mathrm{h}}$ influenced the absolute values of $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ voltages, the device’s sensitivity remained notably consistent. Our analytical model supported these findings and revealed a latching effect with increasing Joule heating. These results suggest that the beam overheat ratio may not be as pivotal as assumed for most thermal MEMS sensors, promising substantial energy consumption reduction.
Third, we investigate the influence of the flow angle on the beam’s performance. We conclude that the device performs similarly to previous hot-wire-based microsensors and can measure flow direction.
We analyzed in detail the beam’s response to fluctuations at different frequencies, typically present in various applications. Our results confirmed that this device, operating near critical transition points, is also sensitive to turbulent flow velocity fluctuations.
Finally, we obtained the first direct measurements of the beam’s shape during transitions and confirmed the theoretical predictions of asymmetric, higher-than-first-mode transitions. We can confirm that the beam’s single-degree-of-freedom approximation needs further improvement.
Overall, this study lays the groundwork for designing a series of flow sensors and multisensor arrays for velocity and velocity gradient measurements using initially curved bistable microbeams. However, unanswered questions still require further investigation, such as the sensor’s response to out-of-plane flow directions, its performance in realistic environments, and the noise levels at which it can accurately measure low- and high-flow velocities for different applications.
@@ Acknowledgments
The authors acknowledge the financial support of the Israeli Innovation Authority, NOFAR program, and RAFAEL LTD, Advanced Defense Systems. S. Krylov acknowledges the support of the Henry and Dinah Krongold Chair of Microelectronics. The authors would also like to thank Gal Spaer Milo, Grigori Gulitski and Stella Lulinsky for their assistance.